# He awa whiria: the tidal streams of interstellar objects

### v1

@@ 4.2 Dynamical Heating
Substantial uncertainty in the appropriate values to use in the heating rate (Ida et al., 1993), and the appropriateness of assuming constant $D$ (Aumer et al., 2016) mean giant molecular clouds may still be an important source of dynamical heating . There are also other sources of heating, including perturbations from spiral arms, the bar, and even cold dark matter substructures (Bovy et al., 2017; Carlberg & Agler, 2023), in rough order of likely importance [^19] Cold dark matter substructures play a much larger role for observed stellar streams than for ISO streams, because the former are on halo-like orbits and hence not subject to the heating mechanisms at play in the disk.. Aumer et al. (2016) argue that molecular clouds likely play an important role in increasing the vertical velocity dispersion, whereas spiral arms primarily increase the velocity dispersion only in the plane of the Galaxy, and these two heating mechanisms evolve in their relative importance over the history of the Galaxy. Moreover, they argue that age uncertainties may artificially lower the power law index of $\sigma$’s variation with $t$, as stars at intermediate ages are incorrectly assigned to both older and younger populations. On top of uncertainties around the heating rates, it is also possible for older stars to be born with higher velocity dispersions (Bournaud et al., 2009; Forbes et al., 2012), driven by a combination of more intense star formation and feedback and greater degrees of gravitational instability at high redshift (e.g. Genzel et al., 2014; Krumholz et al., 2018; Forbes, 2023). Given these complications, we stick to a simple empirical relationship that yields a realistically-large velocity dispersion for the default values, leaving the power law scaling with $t$ equivalent to our free parameter $\alpha_{H}$. Reasonable values of $\mathcal{H}$ range from $\sim 0.1$ (corresponding to heating from GMCs only) to $\sim 1$ (the full AVR is explained by heating of some sort), and $\alpha_{H}$ is likely to be between $0.25$ (GMC heating only) and $0.5$ (constant heating).
@@
Besides the occurrence of multiple ISOs from the same stream, another consequence of the inhomogeneity of ISOs is the time-dependence of the encounter rate as the Sun passes through different streams. In the ISO context this was explored by Portegies Zwart (2021), and it also comes up in the context of dark matter mini-halos (O’Hare et al., 2024). Given that we have a rate distribution that extends to very large values of $\mathcal{R}_{\mathrm{stream}}$, it is worth considering how quickly the Sun passes through individual streams and is able to effectively sample the extremes of the rate distribution. The Sun will transit through a stream of width $\sigma/\kappa$ on a timescale $\sigma\ \kappa^{-1}v_{\mathrm{rel}}^{-1}$, where $v_{\mathrm{rel}}$ is the relative velocity of the Sun and the stream, which will be of order 30 $\mathrm{km}\ \mathrm{s}^{-1}$ for most streams. The resulting timescale is $25\ \mathrm{Myr}(\sigma/v_{\mathrm{rel}})$. Typical values of the latter factor are $\sim 1$, though distributional effects are likely to be important, since streams with higher rates also have lower values of $\sigma$. The Sun has existed for $\sim 180$ such timescales, so streams with rates up to $\mathcal{R}_{\mathrm{stream}}$ such that $N_{\mathrm{stream}}(\mathcal{R}_{\mathrm{stream}})\approx 10^{-2}$ are likely to have interacted with the Sun in its history. Note that these values are again all relative to our normalisation of the encounter rate of ISOs to 10 yr-1. For many of our population models, this $\mathcal{R}_{\mathrm{stream}}$ is of order 10 yr-1, i.e. a single stream producing of order as many ISOs as the currently-expected rate from all streams. This is only a factor-of-two level fluctuation — so intense “showers” of interstellar objects, analogous perhaps in effect but not in cause to the concept of Oort comet showers (e.g. Hills, 1981), are unlikely in Earth’s past or future. Even for a value of $N_{\mathrm{stream}}(\mathcal{R}_{\mathrm{stream}})$ a few orders of magnitude smaller to accommodate distributional effects, the rate distribution is so steep here that $\mathcal{R}_{\mathrm{stream}}$ appears unlikely to exceed $\sim 100\ \mathrm{yr}^{-1}$.
@@ 5 Conclusions
  • The Sun is currently contained in $\sim 10^{7}$ streams from clusters, many of which may be composed of streams from individual stars, totalling $\gtrsim 10^{10}$ streams. However, these streams contribute wildly different ISO encounter rates to the observable volume, ranging from $10^{-15}$ per year, to $\sim 0.1$ per year. The Sun may occasionally pass through streams with far higher rates.
@@
We thank Ana Bonaca, Susanne Pfalzner, Chris Gordon, Joe Masiero, Mordecai-Mark Mac Low, Ciaran O’Hare, Hagai Perets, Rosemary Mardling, Garrett Levine, Darryl Seligman, Jay McMahon, Jack Patterson, and Rua Murray for helpful discussions and comments.
@@ Appendix A Adaptive Anisotropic Kernel Density Estimator
To alleviate this difficulty, we implement an adaptive kernel density estimator (KDE) where the kernel varies from point to point. We still use a multivariate Gaussian kernel centered on $b$, but we set the covariance matrix equal to a multiple $\eta_{\mathrm{cov}}$ of the sample covariance $\Sigma_{\mathrm{samp}}$ of the $N_{\mathrm{neigh}}$ nearest neighbors. This choice increases the importance of the scaling of each parameter, since now the distance between points not only matters in the evaluation of $K_{i}$, but also in determining exactly which points are the $N_{\mathrm{neigh}}$ nearest neighbors, and hence the covariance matrix of each $K_{i}$. We also need to specify the value of $N_{\mathrm{neigh}}$, with smaller numbers being more adaptive to the local density, but also more subject to randomness in the tracers. Regardless, we build a kd tree using scipy.spatial.KDTree (Maneewongvatana & Mount, 1999) when the KDE is initialized to efficiently find the nearest neighbors of each point.
@@
We set the values of the KDE’s hyperparameters ($N_{\mathrm{neigh}}$ and $\eta_{\mathrm{cov}}$) via 5-fold cross-validation. In principle the scaling of each dimension could also be set via cross-validation, but that dramatically increases the dimensionality of the problem. We therefore instead choose physically-motivated scales and only cross-validate 2 hyperparameters. In particular, we set $f_{x}=f_{y}=1$, $f_{z}=\nu(r_{0})/\kappa(r_{0})$, $f_{vx}=f_{vy}=\kappa(r_{0})$, and $f_{vz}=\nu(r_{0})$, where $r_{0}$ is a representative galactocentric radius of the stream, which we set to 8.1 kpc for streams near the Solar neighborhood. To carry out the cross-validation, the dataset is divided at random into $N_{k}=5$ equally-sized subsets. Each of these is, in turn, excluded from the remaining $N_{k}-1=4$ subsets. The 4 subsets are merged into a single training set, which is used to produce a KDE via Equation A1. The excluded subset is then used to test this KDE - its cross-validation error is estimated following Sain (2001) via

### v2

@@ 4.2 Dynamical Heating
Substantial uncertainty in the appropriate values to use in the heating rate (Ida et al., 1993), and the appropriateness of assuming constant $D$ (Aumer et al., 2016) mean giant molecular clouds may still be an important source of dynamical heating . There are also other sources of heating, including perturbations from spiral arms, the bar, and even cold dark matter substructures (Bovy et al., 2017; Carlberg & Agler, 2023), in rough order of likely importance [^19] Cold dark matter substructures play a much larger role for observed stellar streams than for ISO streams, because the former are on halo-like orbits and hence not subject to the heating mechanisms at play in the disk.. Aumer et al. (2016) argue that molecular clouds likely play an important role in increasing the vertical velocity dispersion, whereas spiral arms primarily increase the velocity dispersion only in the plane of the Galaxy, and these two heating mechanisms evolve in their relative importance over the history of the Galaxy (see also Arunima et al. (2025)). Moreover, they argue that age uncertainties may artificially lower the power law index of $\sigma$’s variation with $t$, as stars at intermediate ages are incorrectly assigned to both older and younger populations. On top of uncertainties around the heating rates, it is also possible for older stars to be born with higher velocity dispersions (Bournaud et al., 2009; Forbes et al., 2012), driven by a combination of more intense star formation and feedback and greater degrees of gravitational instability at high redshift (e.g. Genzel et al., 2014; Krumholz et al., 2018; Forbes, 2023). Given these complications, we stick to a simple empirical relationship that yields a realistically-large velocity dispersion for the default values, leaving the power law scaling with $t$ equivalent to our free parameter $\alpha_{H}$. Reasonable values of $\mathcal{H}$ range from $\sim 0.1$ (corresponding to heating from GMCs only) to $\sim 1$ (the full AVR is explained by heating of some sort), and $\alpha_{H}$ is likely to be between $0.25$ (GMC heating only) and $0.5$ (constant heating).
@@
Besides the occurrence of multiple ISOs from the same stream, another consequence of the inhomogeneity of ISOs is the time-dependence of the encounter rate as the Sun passes through different streams. In the ISO context this was explored by Portegies Zwart (2021), and it also comes up in the context of dark matter mini-halos (O’Hare et al. 2024; DSouza et al. (2024)). Given that we have a rate distribution that extends to very large values of $\mathcal{R}_{\mathrm{stream}}$, it is worth considering how quickly the Sun passes through individual streams and is able to effectively sample the extremes of the rate distribution. The Sun will transit through a stream of width $\sigma/\kappa$ on a timescale $\sigma\ \kappa^{-1}v_{\mathrm{rel}}^{-1}$, where $v_{\mathrm{rel}}$ is the relative velocity of the Sun and the stream, which will be of order 30 $\mathrm{km}\ \mathrm{s}^{-1}$ for most streams. The resulting timescale is $25\ \mathrm{Myr}(\sigma/v_{\mathrm{rel}})$. Typical values of the latter factor are $\sim 1$, though distributional effects are likely to be important, since streams with higher rates also have lower values of $\sigma$. The Sun has existed for $\sim 180$ such timescales, so streams with rates up to $\mathcal{R}_{\mathrm{stream}}$ such that $N_{\mathrm{stream}}(\mathcal{R}_{\mathrm{stream}})\approx 10^{-2}$ are likely to have interacted with the Sun in its history. Note that these values are again all relative to our normalisation of the encounter rate of ISOs to 10 yr-1. For many of our population models, this $\mathcal{R}_{\mathrm{stream}}$ is of order 10 yr-1, i.e. a single stream producing of order as many ISOs as the currently-expected rate from all streams. This is only a factor-of-two level fluctuation — so intense “showers” of interstellar objects, analogous perhaps in effect but not in cause to the concept of Oort comet showers (e.g. Hills, 1981), are unlikely in Earth’s past or future. Even for a value of $N_{\mathrm{stream}}(\mathcal{R}_{\mathrm{stream}})$ a few orders of magnitude smaller to accommodate distributional effects, the rate distribution is so steep here that $\mathcal{R}_{\mathrm{stream}}$ appears unlikely to exceed $\sim 100\ \mathrm{yr}^{-1}$.
@@ 5 Conclusions
  • The Sun is currently contained in $\sim 10^{7}$ streams from clusters, many of which may be composed of streams from individual stars, totalling $\gtrsim 10^{10}$ streams. However, these streams contribute wildly different ISO encounter rates to the observable volume, ranging from $10^{-15}$ per year, to $\sim 0.1$ per year. The Sun may occasionally pass through streams with far higher encounter rates.
@@
We thank Ana Bonaca, Susanne Pfalzner, Chris Gordon, Joe Masiero, Mordecai-Mark Mac Low, Ciaran O’Hare, Hagai Perets, Rosemary Mardling, Garrett Levine, Darryl Seligman, Jay McMahon, Jack Patterson, and Rua Murray for helpful discussions and comments. We also appreciate the thoughtful and constructive report from the anonymous referee.
@@ Appendix A Adaptive Anisotropic Kernel Density Estimator
To alleviate this difficulty, we implement an adaptive kernel density estimator (KDE) where the kernel varies from point to point. We still use a multivariate Gaussian kernel centered on $b$, but we set the covariance matrix of the $i$th point equal to a multiple of the sample covariance $\Sigma_{\mathrm{samp}}$ of the $N_{\mathrm{neigh}}$ nearest neighbors:
$\Sigma_{i}=\eta_{\mathrm{cov}}|\Sigma_{i,\mathrm{samp}}|^{\alpha_{\mathrm{cov}}}\Sigma_{i,\mathrm{samp}}$ (A2)
The overall scaling $\eta_{\mathrm{cov}}$ is a constant, and we include another hyperparameter, $\alpha_{\mathrm{cov}}$ to allow the covariance to scale with the determinant of $\Sigma_{i,\mathrm{samp}}$, i.e. with the local density of points. This is a generalization of the adaptive bandwidth estimator proposed by Abramson (1982a) and Abramson (1982b) for which $\alpha_{\mathrm{cov}}=-1/2$ (see Sain, 2002, for a general discussion of multivariate adaptive bandwidths). This choice increases the importance of the scaling of each parameter, since now the distance between points not only matters in the evaluation of $K_{i}$, but also in determining exactly which points are the $N_{\mathrm{neigh}}$ nearest neighbors, and hence the covariance matrix of each $K_{i}$. We also need to specify the value of $N_{\mathrm{neigh}}$, with smaller numbers being more adaptive to the local density, but also more subject to randomness in the tracers. Regardless, we build a kd tree using scipy.spatial.KDTree (Maneewongvatana & Mount, 1999) when the KDE is initialized to efficiently find the nearest neighbors of each point.
@@
We set the values of the KDE’s hyperparameters ($N_{\mathrm{neigh}}$, $\alpha_{\mathrm{cov}}$, and $\eta_{\mathrm{cov}}$) via 5-fold cross-validation. In principle the scaling of each dimension could also be set via cross-validation, but that dramatically increases the dimensionality of the problem. We therefore instead choose physically-motivated scales and only cross-validate 2 hyperparameters. In particular, we set $f_{x}=f_{y}=1$, $f_{z}=\nu(r_{0})/\kappa(r_{0})$, $f_{vx}=f_{vy}=\kappa(r_{0})$, and $f_{vz}=\nu(r_{0})$, where $r_{0}$ is a representative galactocentric radius of the stream, which we set to 8.1 kpc for streams near the Solar neighborhood. To carry out the cross-validation, the dataset is divided at random into $N_{k}=5$ equally-sized subsets. Each of these is, in turn, excluded from the remaining $N_{k}-1=4$ subsets. The 4 subsets are merged into a single training set, which is used to produce a KDE via Equation A1. The excluded subset is then used to test this KDE - its cross-validation error is estimated following Sain (2001) via